OpenStudy (anonymous):

(MEDAL+FAN) FOR BEST ANSWER.(PLEASE HELP) 1.Identify the 35th term of an arithmetic sequence where a1 = -7 and a18 = 95. (2 points) 203 197 168 163 2.Identify the 31st term of an arithmetic sequence where a 1 = 26 and a22 = -226. (2 points) -334 -274 -284 -346

OpenStudy (anonymous):

@gummibearzrule Can u help please

OpenStudy (anonymous):

@mathstudent55 Please Help

OpenStudy (anonymous):

I'm afraid that I'm not extremely experienced at this level, terribly sorry ^^;

OpenStudy (anonymous):

OH Okay Np

OpenStudy (anonymous):

since your first term is -7 you now have your a in the formula: \[a_{n}\] = a + (n-1)d where d is your common difference. does this formula look familiar?

OpenStudy (anonymous):

Yes KINDA

OpenStudy (anonymous):

to find the common differnce set up using this formula: a18 = 95 = -7 + (18-1)d and solve for d

OpenStudy (anonymous):

Okay Let Me Try

OpenStudy (anonymous):

once you have found d then manipulate the equation to solve for the 35th term. this will look like: a35 = -7 + (35-1)d for whatever d turns out to be

OpenStudy (anonymous):

number 2 is going to follow the same rules. does that make more sense?

OpenStudy (anonymous):

Um. So im suppose to solve a18 = 95 = -7 + (18-1)d, I don't understand how to solve it though. Can u walk me through it or help me. PLEASE

OpenStudy (anonymous):

But If I Solve For 5 = -7 + (18-1)d, I Get d=6

OpenStudy (anonymous):

@maddirm96

OpenStudy (anonymous):

so 95= -7 + (18-1)d 95= -7 + 17d 102= 17d 6=d

OpenStudy (anonymous):

now say: a35 = -7 + (35-1)6 = -7 + (34)6 = -7 + 204 =197

OpenStudy (anonymous):

Ohh Okay Thank You. So the Formula for the second one is a31=26+(31-1)d

OpenStudy (anonymous):

almost but not quite. you need to find your common difference before finding your a31 term. start out by solving for d in the equation: a22 = -226 = 26 + (22-1)d just like in the first problem

OpenStudy (anonymous):

So it gives me a22 = -226 = 26 + (22-1)-12 ?

OpenStudy (anonymous):

if -12 became your common difference then you can use that to substitute your variables back into the formula. lol is this sounding familiar at all to you?

OpenStudy (anonymous):

So for every variable i substitute -12 in, except for a

OpenStudy (anonymous):

I Understood The First The First Question. But The Second One...;/

OpenStudy (anonymous):

a is your first term. since the problem gives you your first term as 26 you can use that. but first you need to solve for d. since you got -12 now say: a31 = 26 + (31-1)-12

OpenStudy (anonymous):

So now i solve for 26 + (31-1)-12

OpenStudy (anonymous):

yes! the explicit formula for arithmetic sequences is \[a_{n} = a + (n-1)d\] where n is the term you are trying to find, a is your first term and d is the common difference. \[a_{n} \] will end up being your value at that Nth term

OpenStudy (anonymous):

a31=26+(31-1)-12 =26+(30)-12 =26+-360 =-334

OpenStudy (anonymous):

OMG. I GOT IT THANK YOU SOO MUCH

OpenStudy (anonymous):

@maddirm96

OpenStudy (anonymous):

I MADE A 100.YAY